OpenStudy (anonymous):

Algebra 2 help: Will give a medal

OpenStudy (anonymous):

What do you need?

OpenStudy (anonymous):

find a third degree polynomial equation with rational coefficients that has roots -4 and 6+1 A. x^3-8x^2-11x+148=0 B. x^3-8x^2-12x+37=0 C. x^3-12x^2+37x=0 D. x^3-8x^2-11x=0

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Can you help me with one more question?

OpenStudy (anonymous):

@Jacob194485

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what does descarte's rule of signs tell you about the real roots of the polynomial -2x^3+3x^2-5x-2=0 a. there is one positive root and either 2 or 0 negative roots b. there are either 2 or 0 positive roots and either 2 or 0 negative rootes c. there is one positive root and one negative root d.there are either 2 or 0 positive roots and one negative root

OpenStudy (anonymous):

@jacob194485 that's the question

OpenStudy (anonymous):

d

OpenStudy (anonymous):

hope i helped :)